# Cross-Modal Reconstruction Pretraining for Ramp Flow Prediction at Highway Interchanges

### v1

@@ Abstract
Interchanges are pivotal junctions that facilitate vehicle transfers between highways. Accurate prediction of ramp traffic flow at highway interchanges is essential for proactive traffic management. However, the scarcity of real-time detectors on ramps creates a “real-time blind spot” challenge. While historical data can be obtained through Electronic Toll Collection (ETC) systems, real-time ramp flow cannot be directly used as model input. This data scarcity forces models to rely solely on mainline traffic data to infer complex ramp dynamics. To address this challenge, we propose a novel two-stage framework, the Spatio-Temporal Decoupled Autoencoder (STDAE), which leverages a cross-modal reconstruction pretraining task to mitigate missing data issues. In the pretraining stage, STDAE learns to reconstruct historical ramp flows exclusively from mainline traffic data, thereby compelling the model to capture the intrinsic spatio-temporal relationship between the mainline and ramp traffic flows. Its unique decoupled architecture, consisting of parallel spatial (SAE) and temporal (TAE) autoencoders, efficiently extracts heterogeneous features. In the downstream prediction stage, the learned representations are integrated with specific forecasting models such as GWNet to enhance prediction accuracy. Comprehensive experiments conducted on three real-world interchange datasets demonstrate that our proposed model (STDAEGWNET) consistently outperforms thirteen state-of-the-art baselines across multiple metrics and sampling intervals. Importantly, our method achieves performance comparable to or even surpassing advanced models that directly utilize historical ramp data, thereby validating its effectiveness in overcoming the real-time data gap. The proposed framework provides a robust and practical solution for fine-grained traffic management in detector-scarce scenarios, and its architecture-agnostic nature enables it to serve as a plug-and-play enhancement module to improve diverse forecasting pipelines. Codes and datasets are available at https://github.com/ChaochaoSeu/RPaI-STDAE.
@@
Traffic Flow Prediction , Highway Interchange , Spatial-Temporal Autoencoder , Pretraining , ETC Data
journal: Arxiv
@@ 1 Introduction
Highways are the backbone of transportation networks in urban agglomerations, providing high-capacity mobility for the movement of people and goods between origins and destinations (Guan and He, 2008). The highway network is composed of basic road segments, entrance and exit ramps, and system interchanges (Manual, 2000). As key nodes connecting different highways, interchanges enable traffic flow to achieve continuous and rapid transitions in multiple directions through complex ramp systems. Moreover, interchange structures vary widely, including trumpet, turbine, cloverleaf, and their variants (Garber et al., 1999). However, during peak hours, interchange traffic flow faces severe challenges, as a large number of vehicles performing route changes converge at these nodes, resulting in traffic bottlenecks. These bottlenecks not only reduce interchange efficiency but also propagate along the mainline, ultimately leading to regional congestion (Ma et al., 2024; Mu et al., 2025). Consequently, such congestion seriously undermines both public travel satisfaction (Higgins et al., 2018) and logistics efficiency (Hu et al., 2018). Therefore, to alleviate congestion and maximize highway capacity, it is essential to monitor traffic operations at interchanges and to design advanced and effective management strategies. And accurate ramp traffic flow prediction provides the foundation for refined traffic control (Ma et al., 2020; Liu et al., 2016). Moreover, precise flow prediction helps traffic authorities capture macroscopic demand and analyze the microscopic composition of mainline flow, thereby clarifying the major sources and destinations of vehicles. Furthermore, it supports the design of targeted control strategies and enables dynamic monitoring and management of the transportation network (Wang et al., 2022; Zhang et al., 2022; Han et al., 2020). In addition, ramp traffic flow is indispensable for building complex traffic simulation models and achieving accurate traffic state estimation (Kan et al., 2020).
@@
In practice, however, ramp flow data are frequently unavailable in real time. Not all ramps are equipped with sensors, and some lack communication channels (Nohekhan et al., 2021). Communication failures may also cause partial or complete data loss. For missing ramp detector data, imputation methods can be applied (Duan et al., 2016). Nevertheless, research on estimating ramp flows at locations lacking sensors remains limited. Conventional approaches primarily rely on origin–destination (OD) matrix–based link flow estimation. However, the acquisition of real-time OD matrices is extremely challenging (Marzano et al., 2009). Machine learning (ML) methods also face significant limitations, as they require extensive datasets and are highly dependent on data quality (Shi et al., 2021; Pereira et al., 2022). To overcome these limitations, recent studies have adopted transfer learning frameworks, utilizing mainline traffic flow to estimate ramp flows despite the lack of historical ramp data (Zhang et al., 2023; Ma et al., 2025).
@@
In recent years, with the widespread deployment of Electronic Toll Collection (ETC) gantry systems, high-resolution spatiotemporal traffic data on highways have become increasingly accessible (Tsung et al., 2020; Zhou et al., 2025). Although ETC gantries are rarely installed on ramps due to high costs (Kan et al., 2020), ramp flows can be estimated by matching vehicle records at upstream and downstream gantries (Zhou et al., 2025). However, this approach encounters two primary challenges. First, legal requirements for data privacy and the technical complexity of system integration prevent authorities from processing vehicle records in real time. Consequently, ramps experience a “real-time blind spot”: while historical ramp flows can be obtained for model training, real-time ramp flows are unavailable as input during prediction. Second, data silos among different highway operators impede cross-regional vehicle tracking (Pospisil and Fric, 2016). Even when mainline data are available, equipment failures, adverse weather, or transmission errors often cause missing or anomalous values. Therefore, developing methods to accurately estimate and predict ramp flows from such incomplete mainline data—flows that are inherently difficult to observe—is both a critical technical requirement for fine-grained highway traffic management and a research problem of significant theoretical and practical importance.
@@
Artificial Intelligence (AI) and ML provide promising solutions. From classical time-series models and Support Vector Machines (SVMs) to advanced deep learning models such as Convolutional Neural Networks (CNNs) (Jiang and Luo, 2022), Recurrent Neural Networks (RNNs) (Tedjopurnomo et al., 2020), and Transformers (Wen et al., 2023), these methods have been widely applied to traffic prediction. For ramp flow estimation, studies have used Random Forests (Kan et al., 2020) and deep learning models for hourly-level predictions (Zhang et al., 2023). Their common goal is to capture dynamic spatiotemporal dependencies in traffic data. Nevertheless, ML- and DL-based approaches face several limitations. Most rely on synchronous ramp data as input, which is infeasible under “real-time blind spot” conditions. Many also assume complete datasets, overlooking the prevalence of missing data in real-world scenarios. Transfer learning can mitigate some of these issues but remains limited for complex interchanges. As it does not fully exploit ETC-based historical ramp data. Consequently, existing methods inadequately capture the spatiotemporal coupling between mainline and ramp flows. This highlights the need for a universal model capable of handling diverse interchange structures.
@@
To address these gaps, we propose the Spatio-Temporal Decoupled Autoencoder (STDAE) for ramp flow reconstruction. The framework is inspired by Masked Autoencoders (MAEs) (He et al., 2022), but it introduces a cross-modal reconstruction pretraining task to bridge the ramp “real-time blind spot.” It operates in two stages: (1) mapping learning, where the model reconstructs historical ramp flows from mainline flows, forcing it to learn spatiotemporal mappings; and (2) downstream prediction, where the learned representations are transferred for accurate ramp flow forecasting. The framework requires only real-time mainline data, even if incomplete, making it highly practical. Its design not only addresses missing input data but also improves adaptability and generalization across diverse interchanges.
@@
    Cross-modal reconstruction framework: Uses accessible mainline data to reconstruct ramp flows, addressing the ramp “real-time blind spot.”
@@
    Robust mask-and-predict mechanism: Enhances prediction accuracy under incomplete data conditions.
@@
    Comprehensive real-world validation: Evaluates performance on three real datasets and multiple sampling intervals, demonstrating improved accuracy and cross-scenario applicability.
@@ 2.1 Short-term traffic prediction
Short-term traffic flow prediction is an essential component of Intelligent Transportation Systems (ITS) (Zhao et al., 2019). Accurate traffic flow forecasting provides critical decision support for traffic management authorities and helps effectively prevent and mitigate congestion (Liu et al., 2024b). Early studies mainly relied on classical statistical methods, such as autoregressive integrated moving average (ARIMA) models (Hamed et al., 1995; Zhang, 2003; Williams and Hoel, 2003) and Kalman filter models (Okutani and Stephanedes, 1984). These approaches are structurally simple, easy to implement, and can provide reliable baseline predictions under stable traffic conditions. However, traffic flow is inherently nonlinear and non-stationary, forming a complex time series. Traditional statistical models struggle to capture its intrinsic stochastic fluctuations and complex spatiotemporal dependencies, resulting in limited prediction accuracy under dynamic traffic conditions (Xu et al., 2022).
@@
To overcome these limitations, researchers have increasingly adopted machine learning approaches (Tedjopurnomo et al., 2020). Shallow methods, such as SVM (Chen et al., 2015), K-Nearest Neighbors (KNN) (Zhang et al., 2009), and hybrid models (Li et al., 2023), can flexibly capture nonlinear relationships without requiring strict stationarity assumptions and exhibit strong adaptability to noise and local traffic flow trends. Consequently, these methods achieve higher prediction accuracy under highly fluctuating or complex traffic conditions. With the rapid development of research, deep learning has gradually become the mainstream approach (Tedjopurnomo et al., 2020; Yin et al., 2022; Jiang and Luo, 2022). In particular, RNNs and their variants, especially Long Short-Term Memory (LSTM) networks (Hu et al., 2025) and Gated Recurrent Units (GRU) (Dai et al., 2019), have demonstrated outstanding performance in capturing temporal dependencies. Meanwhile, CNNs have been employed to extract spatial topological features and combined with RNNs to form hybrid models (Jiang and Zhang, 2018; Méndez et al., 2023).
To further characterize the complex topological structure of traffic networks, researchers have proposed spatiotemporal prediction models based on Graph Neural Networks (STGNN) (Wu et al., 2020a; Yu et al., 2018; Zhao et al., 2019; Zheng et al., 2024; He et al., 2025), enhancing the model’s ability to capture spatiotemporal dependencies. In recent years, attention mechanisms and Transformers have also been introduced for traffic prediction (Liu et al., 2024a; Jin et al., 2025), enabling dynamic focus on critical time periods and key road segments, thereby improving the robustness and accuracy of predictions.
@@
It is noteworthy that the majority of these studies have been applied to urban arterials or freeway mainlines where data collection is relatively comprehensive. The development and validation of these methods heavily depend on accessible mainline traffic flow data. In contrast, ramp traffic flows, which are often sparsely instrumented and difficult to observe directly, have received relatively limited attention in the literature, representing a significant gap in current research.
@@ 2.2 Ramp flow prediction
Highway ramp traffic flow prediction is a specialized branch within the field of dynamic traffic flow forecasting. The core task is to estimate the number of vehicles that will enter or exit the freeway mainline via ramps over a given future time interval (Zhang et al., 2023). As noted in Section 1, obtaining ramp flow data is challenging. Thus, existing research has primarily focused on methods for estimating flows at ramps without sensors. Traditional approaches mainly rely on OD matrices to allocate link flows. However, the real-time updating of OD matrices is extremely difficult, making these methods insufficient for dynamic traffic management (Marzano et al., 2009). Some studies combine traffic flow models with OD matrices. But these approaches are typically computationally intensive, exhibit limited accuracy, and require extensive parameter calibration (Chang and Li, 2002).
@@
In recent years, data-driven machine learning and deep learning methods have become mainstream. Nevertheless, a common limitation of these models is their dependence on a large amount of auxiliary information to compensate for missing data. For instance, Nohekhan et al. (Nohekhan et al., 2021) proposed a model that requires static or semi-static inputs such as the annual average daily traffic (AADT), road classification, and free-flow speed. Similarly, Kan et al. (Kan et al., 2020) relied on ramp capacity and peak demand parameters derived from the Highway Capacity Manual (HCM) (Manual, 2000). This strong dependence on auxiliary information not only limits the deployability of models in data-constrained scenarios but also makes them vulnerable to sudden data loss caused by equipment failures or adverse weather conditions.
To alleviate the reliance on training data, recent studies have introduced transfer learning into ramp flow estimation. The core idea of transfer learning is to transfer knowledge learned from a data-rich source domain (e.g., mainline) to a data-scarce target domain (e.g., ramps) (Zhuang et al., 2020), providing a feasible solution for ramps without historical data. Although traffic states differ across highway segments, they exhibit inherent correlations, creating a natural scenario for the application of transfer learning (Zhuang et al., 2020). In ramp flow prediction, transfer learning is mainly applied in two ways: feature-representation transfer, which reduces the discrepancy in marginal distributions between the source and target domains, and parameter transfer, which mitigates differences in conditional distributions. Combining these strategies can significantly improve prediction accuracy and generalization capability in the target domain (Zhang et al., 2023).
However, existing transfer learning methods encounter new challenges when applied to structurally complex interchange scenarios. First, most prior studies focus on ramps with simple structures. Interchanges, however, have diverse layouts and complex ramp systems. This makes it difficult to directly apply existing models. Second, the data environment has evolved. With the deployment of ETC gantries and other technologies, historical ramp flow data can now be obtained through vehicle path matching. As a result, ramps are no longer entirely “data-free.” Historical time-series data are available offline, although real-time data remain inaccessible. This change presents new challenges for conventional methods. Traditional transfer learning methods are designed for fully data-sparse scenarios. They usually rely only on mainline data and ignore the valuable historical ramp data provided by ETC systems. This limits the ability of models to capture the complex spatiotemporal dependencies between mainline and ramp flows. Therefore, the key challenge is to use limited historical data effectively while handling incomplete mainline inputs. This is essential for developing accurate predictive models for ramps in complex interchanges.
2.3 Pretraining framework for traffic prediction
Inspired by successful practices in the field of natural language processing (NLP), pretraining has become a key paradigm for improving model performance (Liu et al., 2023b). The Transformer architecture, due to its strong ability to handle long sequences, forms the core of such pretrained models (Vaswani et al., 2017). In this context, researchers have extended the idea of pretraining to time series modeling. Through self-supervised learning, models extract effective feature representations from large-scale historical data. This enhances the performance of downstream prediction tasks (Shao et al., 2022a; Chi et al., 2025; Gao et al., 2024). Among these methods, MAE (He et al., 2022) has become a mainstream pretraining strategy. The core task of MAE is to mask part of the observed data and reconstruct it using the remaining visible data. This allows the model to learn deep relationships within the data. Such a property makes MAE naturally suitable for handling missing data, providing a new approach to improving the robustness of traffic prediction models.
Existing pretraining frameworks have explored various model structures and masking strategies. Some studies directly apply the MAE paradigm to spatiotemporal data. For example, TST (Zerveas et al., 2021) uses a masked Transformer architecture to learn representations of multivariate time series. Other models, including STEP (Shao et al., 2022a), DGCN-PTL (Chi et al., 2025), and STDMAE (Gao et al., 2024), further integrate the intermediate representations generated by MAE into downstream predictors to improve performance. Notably, STDMAE proposes a spatial-temporal decoupled masking mechanism. It performs masking and reconstruction separately along the temporal and spatial dimensions. This allows the model to better capture spatiotemporal heterogeneity. To model temporal dependencies more precisely, TS-MAE (Liu et al., 2025) constructs a masked autoencoder based on neural ordinary differential equations (Neural ODEs). This enables continuous modeling of evolution patterns. TS-MAE also uses a stepwise masking strategy to enhance learning robustness. In addition, some studies incorporate graph structures into pretraining. PesNet (Fang et al., 2024) uses graph convolution operators to recover masked data. JointSTNet (Cai et al., 2024) introduces a spatial graph capsule module and a temporal gating module. It also combines an adaptive masking strategy to improve the extraction and learning of complex spatiotemporal relationships.
Although these studies demonstrate the advantages of pretraining in capturing long-term and diverse spatiotemporal features, most pretraining tasks focus on self-reconstruction using historical data. That is, they reconstruct missing parts of the same source traffic flow. This “same-modality” reconstruction paradigm works well for general prediction or imputation tasks. However, it cannot directly solve the problem in interchange scenarios. In these cases, data from one modality (mainline traffic) must be used to predict another modality (ramp traffic). To address this, we propose a new cross-modal reconstruction pretraining task. The model uses mainline traffic flow data to reconstruct ramp traffic flows. In this framework, the model can learn more intrinsic patterns between mainline and ramp flows. Moreover, masking can be applied to the mainline inputs. This simulates real-world scenarios with partial missing data and enhances the robustness of the prediction model.
In summary, although ramp flow prediction makes great progress, current models still have some limitations. To compare the strengths and weaknesses of different approaches more clearly, this study reviews ramp flow prediction methods together with recent pretraining frameworks. Traditional approaches mainly use traffic flow models and dynamic OD estimation. They can predict ramp flows under certain conditions. However, they require complete and real-time data, and the computation cost is high. Machine learning methods reduce the impact of missing data by using additional static or semi-static features. But they rely heavily on auxiliary information, which limits scalability. Transfer learning methods transfer knowledge from mainline flows to ramps. This improves prediction accuracy, but they remain insufficient in complex interchange scenarios. Pretraining approaches take a different path. They learn representations from large-scale historical data through self-supervised learning. These methods show strong robustness and generalization. They also perform well under missing data conditions.
@@ 2.3 Pretraining framework for traffic prediction
Model1 Target2 Data Conditions3 Modeling Paradigm4 Model Evaluation5 Reference IDS DI DT Type TDM SDM VM RTC SGA OD RFP M-S, R-S M-C, R-C RT T $\checkmark$ $\times$ SB $\checkmark$ L Chang and Li (2002) RF/GBM RFP M-S M-C, R-M RT ML $\checkmark$ $\times$ RW $\checkmark$ M Kan et al. (2020) DDA+MT RFP M-S M-C, R-M RT TL $\checkmark$ $\times$ RW $\checkmark$ M Zhang et al. (2023) TST MFP M-S R Pre-N, Test-RT Pre $\checkmark$ $\times$ RW $\checkmark$ H Zerveas et al. (2021) STEP NFP M-LS R Pre-N, Test-RT Pre $\checkmark$ $\checkmark$ RW $\checkmark$ H Shao et al. (2022a) DGCN-PTL NFP M-LS R Pre-N, Test-RT Pre $\checkmark$ $\checkmark$ RW $\checkmark$ H Chi et al. (2025) STDMAE NFP M-LS R Pre-N, Test-RT Pre $\checkmark$ $\checkmark$ RW $\checkmark$ H Gao et al. (2024) TS-MAE MFP M-S R Pre-N, Test-RT Pre $\checkmark$ $\times$ RW $\checkmark$ H Liu et al. (2025) PreSTNet NFP M-LS R Pre-N, Test-RT Pre $\checkmark$ $\checkmark$ RW $\checkmark$ H Fang et al. (2024) JointSTNet NFP M-S R Pre-N, Test-RT Pre $\checkmark$ $\checkmark$ RW $\checkmark$ H Cai et al. (2024)
Note: 1 Model: Name of the traffic flow prediction model evaluated.
@@
2 Target: Application domain from which the dataset is derived, including Ramp Flow Prediction (RFP), Mainline Flow Prediction (MFP), or Network-Wide Flow (NFP).
@@
3 Data Conditions: Assessment of the dataset used, including Input Data Source (IDS), Data Integrity (DI), and Data Timeliness (DT).
@@
4 Modeling Paradigm: Type of model employed, including Traffic Flow Model (T), Machine Learning (ML), Transfer Learning (TL), or pretraining (Pre), along with Temporal Dependency Modeling (TDM) and Spatial Dependency Modeling (SDM) capabilities.

### v2

# Proxy Reconstruction Pre-training for Ramp Flow Prediction at Highway Interchanges

@@ Abstract
Interchanges are junctions where traffic merges, diverges and weaves between highways. Accurate prediction of ramp traffic flow at highway interchanges is essential for proactive traffic management involving flow control and ramp management. In practice, real-time mainline flow can be obtained through Electronic Toll Collection (ETC) systems and ramp flow can be inferred from license plate matching. However, data privacy restrictions and matching latency prevent real-time access to ramp data, leading to a ”real-time blind spot”. This discrepancy dictates that while the model can utilize both mainline and ramp flow data during the training phase, the input is strictly limited to the real-time observable mainline flow during actual deployment. To address this challenge, this paper propose a novel two-stage framework, including pre-training and prediction. In the pre-training stage, the Spatio-Temporal Decoupled Autoencoder (STDAE) leverages a proxy reconstruction task to mitigate missing ramp data issues. STDAE learns to reconstruct historical ramp flows exclusively from mainline traffic data, thereby compelling the model to capture the intrinsic spatio-temporal relationship between the mainline and ramp traffic flows. The unique decoupled architecture of STDAE, consisting of parallel spatial (SAE) and temporal (TAE) autoencoders, efficiently extracts the temporal and spatial dependency features of traffic flow. In the downstream prediction stage, the learned representations are integrated with specific forecasting models such as GWNet to enhance prediction accuracy. Comprehensive experiments conducted on three real-world interchange datasets demonstrate that our proposed model (STDAEGWNET) consistently outperforms thirteen state-of-the-art baselines across multiple metrics and sampling intervals. The proposed method attains performance comparable to, or even exceeding, that of advanced models utilizing historical ramp data. These findings confirm the capability to effectively mitigate the real-time data gap. Furthermore, our model demonstrates strong robustness against missing ETC data scenarios. Importantly, the architecture-agnostic nature enables STDAE to serve as a plug-and-play enhancement module to improve diverse forecasting pipelines.
@@
Traffic Flow Prediction , Highway Interchange , Spatial-Temporal Autoencoder , Pre-training , ETC Data
journal: Applied Soft Computing
@@ 1 Introduction
Highways are the backbone of transportation networks in urban agglomerations, providing high capacity for the movement of people and goods between origins and destinations [guan2008statistical]. The highway network is composed of basic road segments, entrance and exit ramps, and system interchanges [manual2000highway]. As key nodes connecting different highways, interchanges enable traffic flow to achieve continuous and rapid transitions in multiple directions through complex ramp systems. Interchange structures vary widely, including trumpet, turbine, cloverleaf, and their variants [garber1999guidelines]. During peak hours, a large number of vehicles performing route changes converge at these nodes, resulting in traffic bottlenecks. These bottlenecks reduce interchange efficiency and propagate along the mainline, ultimately leading to regional congestion [ma2024analysis, mu2025compliance]. Such congestion seriously undermines public travel satisfaction [higgins2018all] and logistics efficiency [hu2018joint]. Therefore, it is essential to monitor traffic operations at interchanges and to design advanced and effective management strategies. Accurate ramp traffic flow prediction provides the foundation for refined management [ma2020statistical, liu2016model]. It helps traffic authorities analyze the of mainline flow composition, identify sources and destinations, and design targeted control strategies for dynamic network management [wang2022make, zhang2022adapgl, han2020hierarchical]. Ramp flow data also support traffic simulation and accurate state estimation [kan2020novel].
Figure 1: Illustration of the three primary challenges faced by ramp flow prediction based on mainline data. (a) Data silos: vehicle records are fragmented among different highway operators, preventing cross-regional vehicle tracking. (b) Data loss: sensor failures, adverse weather, or transmission errors can cause missing values even in mainline data. (c) Real-time blind spot: real-time ramp flow data are unavailable due to privacy restrictions and system integration barriers, although historical data can be used for model training.
@@
Ramp flow data are frequently unavailable because many ramps lack sensors or communication channels [nohekhan2021deep]. For missing ramp detector data, imputation methods can be applied [duan2016efficient]. Nevertheless, research on estimating ramp flows at locations lacking sensors remains limited. Conventional approaches primarily rely on the estimation of origin–destination (OD) matrix flows. However, the acquisition of real-time OD matrix is extremely challenging [marzano2009limits]. Machine learning (ML) methods also face significant limitations, as they require extensive datasets and are highly dependent on data quality [shi2021physics, pereira2022short]. To overcome these limitations, recent studies have adopted transfer learning frameworks, utilizing mainline traffic flow to estimate ramp flows despite the lack of ramp flow data [zhang2023transfer, ma2025data].
@@
With the widespread deployment of Electronic Toll Collection (ETC) gantry systems, high-resolution spatiotemporal traffic data on highways become increasingly accessible [tsung2020visualizing, zhou2025feasibility]. Although ETC gantries are rarely installed on ramps due to high costs [kan2020novel], ramp flows can be estimated by matching vehicle records at upstream and downstream gantries [zhou2025feasibility]. However, this approach encounters two primary challenges, as illustrated in Figure 1. First, legal requirements for data privacy and the technical complexity of system integration prevent authorities from processing vehicle records in real time. Consequently, ramps experience a “real-time blind spot”: while historical ramp flows can be obtained for model training, real-time ramp flows are unavailable as input during prediction. Second, data silos among different highway operators impede cross-regional vehicle tracking [pospisil2016implementation]. Even when mainline data are available, equipment failures, adverse weather, or transmission errors often cause missing values. Therefore, developing a method that can accurately predict ramp flows based on mainline data, while also accounting for potential data loss, is a critical technical requirement for refined highway traffic management and a research topic of significant theoretical value.
@@
Artificial Intelligence (AI) and ML methods are applied to address the challenges of ramp traffic flow prediction, offering excellent capabilities in capturing complex nonlinear relationships and spatiotemporal dependencies. From classical time series models and Support Vector Machines (SVMs) to advanced deep learning models such as Convolutional Neural Networks (CNNs) [jiang2022graph], Recurrent Neural Networks (RNNs) [tedjopurnomo2020survey], and Transformers [ijcai2023p759], these methods are widely applied to traffic prediction. For ramp flow estimation, studies use Random Forests [kan2020novel] and deep learning models for hourly predictions [zhang2023transfer]. Their common goal is to capture dynamic spatiotemporal dependencies in traffic data. Nevertheless, these approaches face limitations as most require synchronous ramp data, which is infeasible under “real-time blind spot” conditions. Many also assume complete datasets, overlooking the prevalence of missing data in real world scenarios. Transfer learning can mitigate some of these issues but remains limited for complex interchanges. It inadequately captures the spatiotemporal coupling between mainline and ramp flows, as it does not fully exploit ETC-based historical ramp data. This highlights the need for a universal model capable of handling diverse interchanges.
@@
To address these gaps, this paper propose the Spatio-Temporal Decoupled Autoencoder (STDAE) for ramp flow reconstruction. The framework is inspired by Masked Autoencoders (MAEs) [he2022masked], but it introduces a proxy reconstruction pre-training task to bridge the ramp “real-time blind spot” and missing data. It operates in two stages: (1) mapping learning, where the model reconstructs historical ramp flows from mainline flows, forcing it to learn spatiotemporal mappings; and (2) downstream prediction, where the learned representations are transferred for accurate ramp flow forecasting. The framework requires only real-time mainline data when deployed, even if incomplete, making it highly practical. Its design not only addresses missing input data but also improves adaptability and generalization across diverse interchanges.
@@
    A proxy reconstruction framework for ramp flow data: This framework reconstructs ramp flow data, which is difficult to obtain, using simpler and more easily observable mainline flow sequences. This effectively addresses the model deployment challenges caused by ramp ”real-time data blind spot”.
@@
    Prediction task design for missing data: Drawing on the principles of MAE, this framework designs temporal and spatial masks before encoding to simulate missing data. This allows the model to reconstruct complete ramp data from missing mainline data, thereby improving the accuracy of the prediction model.
@@
    Comprehensive real-world validation: The reconstruction and prediction performance is systematically evaluated on three real highway interchange datasets. A commonly used sampling interval of 5 min is selected [WOS:001347142804014, shao2024exploring], along with 3 min and 10 min, to assess the model’s adaptability to different temporal requirements. The results demonstrate the effectiveness of the proposed framework.
@@ 2.1 Short-term traffic prediction
Short-term traffic flow prediction is an essential component of Intelligent Transportation Systems (ITS) [zhao2019t]. Accurate traffic flow forecasting provides critical decision support for traffic management authorities and helps effectively prevent and mitigate congestion [liu2024spatial]. Early studies mainly relied on classical statistical methods, such as autoregressive integrated moving average (ARIMA) models [hamed1995short, zhang2003time, williams2003modeling] and Kalman filter models [okutani1984dynamic]. These approaches are structurally simple, easy to implement, and can provide reliable baseline predictions under stable traffic conditions. However, traffic flow is inherently nonlinear and non-stationary, forming a complex time series. Traditional statistical models struggle to capture its intrinsic stochastic fluctuations and complex spatiotemporal dependencies, resulting in limited prediction accuracy under dynamic traffic conditions [9790148].
@@
Researchers increasingly adopt machine learning approaches to capture nonlinear relationships in traffic flow [tedjopurnomo2020survey]. Shallow methods, such as SVM [chen2015forecasting], K-Nearest Neighbors (KNN) [zhang2009short], and hybrid models [li2023qpso], can flexibly capture nonlinear relationships without requiring strict stationarity assumptions and exhibit strong adaptability to noise and local traffic flow trends. Consequently, these methods achieve higher prediction accuracy under complex traffic conditions. With the rapid development of research, deep learning has gradually become the mainstream approach [tedjopurnomo2020survey, 9352246, jiang2022graph]. In particular, RNNs and their variants, especially Long Short-Term Memory (LSTM) networks [hu2025novel, dong2025real] and Gated Recurrent Units (GRU) [dai2019short], demonstrate outstanding performance in capturing temporal dependencies. Meanwhile, CNNs are employed to extract spatial topological features and combined with RNNs to form hybrid models [jiang2018geospatial, mendez2023long]. To further characterize the complex topological structure of traffic networks, researchers propose spatiotemporal prediction models based on Graph Neural Networks (STGNN) [wu2020comprehensive, WOS:000764175403107, zhao2019t, 10151673, HE2025127564], enhancing the model’s ability to capture spatiotemporal dependencies. Aattention mechanisms and Transformers are also introduced for traffic prediction [liu2024itransformer, jin2025ada, dong2025transformer], enabling dynamic focus on critical time periods and key road segments, thereby improving the robustness and accuracy of predictions.
@@
It is noteworthy that the majority of these studies are applied to urban arterials or freeway mainlines where data collection is relatively comprehensive. The development and validation of these methods heavily depend on accessible mainline traffic flow data. In contrast, ramp traffic flows, which are often sparsely instrumented and difficult to observe directly, receive relatively limited attention in the literature, representing a significant gap in current research.
@@ 2.2 Ramp flow prediction
Highway ramp traffic flow prediction is a specialized branch within the field of traffic flow forecasting. The core task is to estimate the number of vehicles that will merge and diverge via ramps over a given future time interval [zhang2023transfer]. As noted in Section 1, obtaining ramp flow data is challenging. Thus, existing research has primarily focused on methods for estimating flows at ramps without sensors. The first three rows of Table 1 provide a comprehensive comparison of different types of ramp traffic flow prediction models. Traditional approaches mainly rely on OD matrices to allocate link flows. However, the real-time updating of OD matrices is extremely difficult (RTC: ✗), making these methods insufficient for dynamic traffic management [marzano2009limits]. Some studies combine traffic flow models with the OD matrix, but these methods typically require complete mainline and ramp data (DI: M-C, R-C) and are computationally intensive [chang2002optimization]. They also struggle to cope with missing data.
@@
Data-driven methods have become mainstream due to their powerful feature capture capabilities. As shown in Table 1, this strong reliance on auxiliary information limits the model’s scenario generalization ability (SGA). For instance, Nohekhan et al. [nohekhan2021deep] proposed a model that requires static or semi-static inputs such as the annual average daily traffic (AADT), road classification, and free-flow speed. Similarly, Kan et al. [kan2020novel] relied on ramp capacity and peak demand parameters derived from the Highway Capacity Manual (HCM) [manual2000highway]. The model requires the integrity of the mainline data (DI: M-C), making it susceptible to data loss.
@@
Model1 Target2 Data Conditions3 Modeling Paradigm4 Model Evaluation5 Reference IDS DI DT Type TDM SDM VM RTC SGA OD RFP M-S, R-S M-C, R-C RT T $\checkmark$ $\times$ SB L [chang2002optimization] RF/GBM RFP M-S M-C, R-M RT ML $\checkmark$ $\times$ RW $\checkmark$ M [kan2020novel] DDA+MT RFP M-S M-C, R-M RT TL $\checkmark$ $\times$ RW $\checkmark$ M [zhang2023transfer] TST MFP M-S R Pre-N, Test-RT Pre $\checkmark$ $\times$ RW $\checkmark$ H [zerveas2021transformer] STEP NFP M-LS R Pre-N, Test-RT Pre $\checkmark$ $\checkmark$ RW $\checkmark$ H [shao2022pre] DGCN-PTL NFP M-LS R Pre-N, Test-RT Pre $\checkmark$ $\checkmark$ RW $\checkmark$ H [chi2025dynamic] STDMAE NFP M-LS R Pre-N, Test-RT Pre $\checkmark$ $\checkmark$ RW $\checkmark$ H [WOS:001347142804014] TS-MAE MFP M-S R Pre-N, Test-RT Pre $\checkmark$ $\times$ RW $\checkmark$ H [liu2025ts] PreSTNet NFP M-LS R Pre-N, Test-RT Pre $\checkmark$ $\checkmark$ RW $\checkmark$ H [fang2024prestnet] JointSTNet NFP M-S R Pre-N, Test-RT Pre $\checkmark$ $\checkmark$ RW $\checkmark$ H [cai2024jointstnet] Ours RFP M-LS, R-LS R Pre-N, Test-RT Pre $\checkmark$ $\checkmark$ RW $\checkmark$ H
Note: 1 Model: Name of the evaluated traffic flow prediction model.
@@
2 Target: Application domain, including Ramp Flow Prediction (RFP), Mainline Flow Prediction (MFP), and Network-Wide Flow Prediction (NFP).
@@
3 Data Conditions: Dataset assessment including Input Data Source (IDS), Data Integrity (DI), and Data Timeliness (DT); M-S/M-LS: Mainline Short/Long–Short series; R-S/R-LS: Ramp Short/Long–Short series; M-C/R-C: Mainline/Ramp Complete data; R-M: Ramp Missing data; R: Robust to missing data.
@@
4 Modeling Paradigm: Includes model type (Type), Temporal Dependency Modeling (TDM), and Spatial Dependency Modeling (SDM); T: Traffic Flow Model, ML: Machine Learning, TL: Transfer Learning, Pre: Pre-training Model.